Simplify
2- 1/3
--------
4 + 11/3

Answer by sdmmadam@yahoo.com(530)

Simplify
2- 1/3
--------
4 + 11/3
= (2- 1/3) 3
---------- X ----
(4 + 11/3) 3
= (2- 1/3)X3
------------
(4 + 11/3)X3
= [(2X3)- (1/3)X3]
---------------
[(4X3) + (11/3)X3]
=(6-1)/(12+11)
=5/23